# Helical randomization of magnetized Galactic plasmas: from magnetorotational disc dynamo to the Faraday rotation sky

### v1

@@ 5 Conclusions
The above-considered results obtained in numerical simulations (models) and inferred from the observations of the Faraday rotation sky show that magnetic helicity dominates the chaotic/turbulent dynamics (directly or through the Kolmogorov-Iroshnikov phenomenology) in the internal accretion disk plasma around the central black hole and the global Galactic plasma’s magnetic field and electron density. The magnetic field imposes its level of randomization on the plasma’s electron density .
This helcal randomization takes place even in the case of zero (or negligible) net helicity due to the phenomenon of spontaneous breaking of the local reflectional symmetry build-in the chaotic/turbulent plasma’s dynamics.
The applicability of the above-suggested approach to the kinetic scales (that is relevant to the internal disk’s collisionless plasma) has been also briefly discussed
Despite very considerable differences in the physical parameters and scales the results of the numerical simulations (models) related to the randomization phenomenon are in quantitative agreement with the observations of the Galactic plasmas.

### v4

@@ 7 Conclusions
The above-considered results obtained in numerical simulations (models) and inferred from the observations of the Faraday rotation and synchrotron emission maps show that magnetic helicity dominates the chaotic/turbulent dynamics (directly or through the Kolmogorov phenomenology) in the internal accretion disk plasma around the central black hole and in the global Galactic plasma’s magnetic field, electron density, and the above-mentioned maps. The magnetic field imposes its level of randomization on the plasma’s electron density and the Faraday rotation and synchrotron emission maps. Analogous conclusions can also be true for galaxy intracluster plasmas.
The dominated by magnetic helicity randomization occurs even in the case of zero (or negligible) net helicity due to the spontaneous breaking of the local reflectional symmetry (intrinsic to the chaotic/turbulent plasma’s dynamics).
The applicability of the above-suggested approach to the kinetic scales (relevant to the internal Galactic disk’s collisionless plasma) has also been established.
Despite considerable differences in the physical parameters and scales, the results of the numerical simulations (models) related to the randomization phenomenon are in quantitative agreement with the observations of the plasmas of galaxies and galaxy clusters.